OpenStudy (anonymous):

\[\int_{-\infty}^1 e^x dx\] =\[e^x\bigg|_{-\infty}^{1}\] I know....

OpenStudy (lgbasallote):

improper integral huh

OpenStudy (lgbasallote):

hint \[e^{-\infty} = 0\]

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

yep...but the set up is ok?

OpenStudy (lgbasallote):

yes @MathSofiya that's right

OpenStudy (anonymous):

I've been separating integrals for the past hour...so that was strange....lol

OpenStudy (anonymous):

I thought I had to do something like \[\lim_{t \rightarrow \infty}\]

OpenStudy (lgbasallote):

well that's the proper way...

OpenStudy (helder_edwin):

actually \[ \Large \int_{-\infty}^1e^x\,dx=\lim_{b\to-\infty}\int_b^1e^x\,dx= \lim_{b\to-\infty}(e^1-e^b) \]

OpenStudy (anonymous):

To be precise, one has to find \[ \lim_{c\to -\infty} \int_c^1 e^x dx =\\ \lim_{c\to -\infty} e^1 - e^c=e- 0=e \]

OpenStudy (lgbasallote):

\[\int_{-\infty}^1 e^x dx\] \[\lim_{t\rightarrow -\infty} \int_t^1 e^xdx\] \[e^x|_t^1\] \[e^1 - e^t\] \[e - e^{-\infty}\] \[e - 0\] \[e\]

OpenStudy (anonymous):

thanks...that's more like it...the whole section was along the lines of \[\lim_{t \rightarrow \infty}\] and this integral seemed too easy Thanks again

OpenStudy (lgbasallote):

as long as you know \(e^{-\infty} = 0\) it's easy

OpenStudy (anonymous):

yes sir!